# Generic consistency and nondegeneracy of vertically parametrized systems

### v3

@@ 1. Introduction
In a follow-up paper, we use the results from this paper to strengthen several previous statements about reaction networks, in particular concerning absolute concentration robustness from [PEF22] and nondegenerate multistationarity from [CFMW17].
@@
We let $\star\colon\mathbb{C}^{n}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ denote the Hadamard product, given by $({x}\star{y})_{i}=x_{i}y_{i}$. For a field $K$, we let $K^{*}=K\setminus\{{0}\}$ be the group of units. For a matrix $A=(a_{ij})\in\mathbb{Z}^{n\times m}$ and a vector ${x}\in(K^{*})^{n}$, we let ${x}^{A}\in K^{m}$ be defined by $({x}^{A})_{j}=x_{1}^{a_{1j}}\cdots x_{n}^{a_{nj}}$ for $j=1,\ldots,m$. The zero matrix of size $n\times m$ is denoted $0_{n\times m}$, and ${\rm Id}_{n}$ is the identity matrix of size $n$. When saying that a property holds generically in a family indexed by some parameters in a set $\mathcal{P}\subseteq\mathbb{C}^{k}$, we mean that it holds in a nonempty Zariski open subset of $\mathcal{P}$. For a set $S\subseteq\mathbb{C}^{n}$, we let $\overline{S}$ denote the Zariski closure of $S$ in $\mathbb{C}^{n}$.
@@ 2.1. Framework
$\pi\colon\mathcal{I}\rightarrow\mathbb{C}^{k},\quad(p,{x})\mapsto p.$
For each choice of parameters $p\in\mathbb{C}^{k}$, we get a specialized system
@@
By Chevalley’s theorem [Sta24, 00F5], $\mathcal{Z}$ is constructible as it is the image of a constructible set under a projection map. Therefore, if $\mathcal{Z}$ is Zariski dense, it must contain a nonempty Zariski open set of $\mathbb{C}^{k}$. In this case, a property holds generically in $\mathcal{Z}$ if and only if it does in $\mathbb{C}^{k}$.
@@
The system $F$ is implicit in the notation.
With this notation in place, we proceed now to study the generic dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ for $p\in\mathcal{P}$ in relation to whether the system is generically consistent, in the sense that $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense in $\mathbb{C}^{k}$. We then move on to study zeros in subsets $\mathcal{X}\subseteq(\mathbb{C}^{*})^{n}$. We also relate the study of nonemptiness and dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ to the concept of nondegeneracy. The connections among these properties are summarized in Theorem 2.15 towards the end of this section.
@@ 2.2. Generic consistency and dimension
Let $F\in\mathbb{C}[p,{x}^{\pm}]^{r}$ and let $\mathcal{P}\subseteq\mathbb{C}^{k}$ be Zariski dense. Assume that the incidence variety $\mathcal{I}$ is irreducible, and let $d:=\dim(\mathcal{I})-k$. The following holds:
@@
    If $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense in $\mathbb{C}^{k}$, then
    $\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))=d$
    for all $p$ in a nonempty Zariski open set $\mathcal{U}\subseteq\mathcal{Z}_{\mathcal{P}}$. Furthermore, if $\mathcal{I}$ has the expected dimension (so that $d=n-r$), then $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ has pure dimension $n-r$ for all $p\in\mathcal{U}$.
@@
    If $\mathcal{Z}_{\mathcal{P}}$ is not Zariski dense in $\mathbb{C}^{k}$, then for all $p\in\mathcal{Z}_{\mathcal{P}}$ it holds that
    $\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))>d.$
@@
We apply the Theorem on the Dimension of Fibers (see, e.g., [Sha13, Theorem 1.25]) to the canonical projection map
$\pi\colon\mathcal{I}\rightarrow\overline{\mathcal{Z}},\quad(p,x)\mapsto p,$
which is a dominant morphism of irreducible varieties, to conclude that for all $p\in\mathcal{Z}$, we have
$\dim(\pi^{-1}(p))\geq\dim(\mathcal{I})-\dim(\,\overline{\mathcal{Z}}\,)=k+d-\dim(\,\overline{\mathcal{Z}}\,)=:e.$ (2.4)
Moreover, there exists a nonempty Zariski open subset $U\subseteq\overline{\mathcal{Z}}$ such that, for any $p\in U$, it holds that $\dim(\pi^{-1}(p))=e$ or $\pi^{-1}(p)=\varnothing$.
Let us consider case (i). As $\mathcal{Z}_{\mathcal{P}}\subseteq\mathcal{Z}$, taking Zariski closures in $\mathbb{C}^{k}$ gives $\mathbb{C}^{k}=\overline{\mathcal{Z}_{\mathcal{P}}}\subseteq\overline{\mathcal{Z}}$, and hence $\overline{\mathcal{Z}}=\mathbb{C}^{k}$. Therefore, $e=d$. It follows that
$\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))=d$
for all $p\in U\cap\mathcal{Z}_{\mathcal{P}}$, since in this case $\pi^{-1}(p)=\mathbb{V}_{\mathbb{C}^{*}}(F_{p})\neq\varnothing$. Noting that $U\cap\mathcal{Z}_{\mathcal{P}}\neq\varnothing$ as $\mathcal{Z}_{\mathcal{P}}\subseteq\mathbb{C}^{k}$ is Zariski dense, we obtain statement (i) by taking $\mathcal{U}=U\cap\mathcal{Z}_{\mathcal{P}}$.
Consider now case (ii). By (2.4), the result follows if $\overline{\mathcal{Z}}\subseteq\mathbb{C}^{k}$ is proper, and hence of dimension smaller than $k$. If on the contrary $\overline{\mathcal{Z}}=\mathbb{C}^{k}$, then $\mathcal{Z}$ contains a nonempty Zariski open subset $U\subseteq\mathbb{C}^{k}$, and hence $U\cap\mathcal{P}\subseteq\mathcal{Z}\cap\mathcal{P}=\mathcal{Z}_{\mathcal{P}}$. Since $\mathcal{P}$ is Zariski dense in $\mathbb{C}^{k}$, we obtain $\mathbb{C}^{k}=\overline{U\cap\mathcal{P}}\subseteq\overline{\mathcal{Z}_{\mathcal{P}}}$, which is a contradiction. This shows that $\overline{\mathcal{Z}}$ is proper and concludes the proof of (ii). ∎
Combining cases (i) and (ii) of Theorem 2.2, we obtain that $\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))=d$ for some $p$ if and only if $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense in $\mathbb{C}^{k}$. Furthermore, if the dimension is $d$ for one value of $p$, then it is generically $d$ in $\mathcal{Z}_{\mathcal{P}}$.
@@
Consider the system in (1.8) where $n=k=r=2$. We have $\mathcal{Z}=(\mathbb{C}^{*})^{2}$. Furthermore, $\mathcal{I}$ is irreducible of dimension $3$. By Theorem 2.2(i), it holds that $\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))=3-2=1$ for generic $p\in(\mathbb{C}^{*})^{2}$, which we already saw in (1.8). In this case, the generic dimension differs from the expected dimension $n-r=0$ from (2.3).
@@ 2.3. Restricting the ambient space
but equality might not hold (consider $F=(x_{1}-1)^{2}+(x_{2}-1)^{2}$ and $\mathcal{X}=\mathbb{R}^{2}_{>0}$). In general, equality in (2.5) holds if $\mathcal{X}$ is a Euclidean open subset of $(\mathbb{C}^{*})^{n}$, or if $\mathcal{X}$ is a Euclidean open subset of $(\mathbb{R}^{*})^{n}$ and additionally each irreducible component of $\mathbb{V}_{\mathbb{C}^{*}}^{\mathcal{X}}(F)$ contains a nonsingular real point (see [PEF22, Theorem 6.5] and [BCR98, Prop. 3.3.16]). Furthermore, if $F\in\mathbb{R}[x^{\pm}]^{r}$, then $n-r$ is a lower bound of the dimension of $\mathbb{V}_{\mathbb{R}^{*}}(F)$ if it is nonempty and each of its irreducible components contains a nonsingular point. Similarly, $n-r$ is a lower bound on the dimension of $\mathbb{V}_{\mathbb{R}^{*}}(F)\cap\mathbb{R}^{n}_{>0}$ as a semialgebraic set if it is nonempty and each irreducible component of $\mathbb{V}_{\mathbb{R}^{*}}(F)$ that intersects >$\mathbb{R}^{n}_{>0}$ contains a nonsingular point.
@@
If $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$ is Zariski dense in $\mathbb{C}^{k}$, then so is $\mathcal{Z}_{\mathcal{P}}$. If $\mathcal{I}$ is irreducible and attains its expected dimension $k+n-r$, then Theorem 2.2(i) readily gives that $\mathbb{V}_{\mathbb{C}^{*}}^{\mathcal{X}}(F_{p})$ has pure dimension $n-r$ for generic $p\in\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$. The following example illustrates that Theorem 2.2(ii) might not extend to subsets $\mathcal{X}$ unless extra requirements are imposed.
@@
In particular $\mathcal{Z}$ is Zariski dense if and only if $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$ is Zariski dense.
@@
Assume that $\mathcal{I}$ is nonsingular and irreducible. If $\mathcal{P}$ and $\mathcal{X}$ are Euclidean open subsets of $\mathbb{C}^{k}$ and $(\mathbb{C}^{*})^{n}$, respectively, then $\mathcal{I}\cap(\mathcal{P}\times\mathcal{X})$ is Zariski dense in $\mathcal{I}$ if this intersection is nonempty. The same conclusion holds if $F$ has real coefficients and $\mathcal{P}$ and $\mathcal{X}$ are Euclidean open subsets of $\mathbb{R}^{k}$ and $\mathbb{R}^{n}$, respectively. This follows from the nonsingularity of $\mathcal{I}$, together with the fact that for a complex irreducible variety defined by polynomials with real coefficients that has at least one real nonsingular point, the real points form a Zariski dense subset of the complex variety [BCR98, Prop. 3.3.16]. Observe that in Example 2.5, the intersection $\mathcal{I}\cap(\mathcal{P}\times\mathcal{X})$, which is not dense in $\mathcal{I}$, consists only of singular points.
@@ 2.4. Nondegeneracy, Euclidean interior, and (real) dimension
Note that, unlike the weaker notion of nonsingularity, nondegeneracy depends on the particular tuple $F\in\mathbb{C}[x^{\pm}]^{r}$, and not just the variety $\mathbb{V}_{\mathbb{C}^{*}}(F)$. The relationship between nondegenerate zeros and nonsingular points, and the connection to dimension, is summarized in the following proposition, which gathers several well-known results.
@@
Statement (i) is [CLO15, Thm 9.6.9]. For (ii), the uniqueness follows from [CLO15, Theorem 9.6.8(iv)], whereas the density and dimensionality claims follow from Remark 2.4 (see also [BCR98, Theorem 3.3.10]). For statement (iii), consider the Jacobian $J_{F}=(\partial f_{i}/\partial x_{j})\in\mathbb{C}[p,x^{\pm}]^{r\times n}$, and note that the complement $\mathcal{I}\setminus\mathcal{U}\subseteq\mathcal{I}$ is cut out by the $r$-minors of $J_{F}$, and therefore forms a Zariski closed subset of $\mathcal{I}$. ∎
Theorem 2.2(i) allows us to determine the generic dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ from the Zariski denseness of $\mathcal{Z}_{\mathcal{P}}$, while the implicit function theorem (as we will use below in Proposition 2.12) gives that nondegeneracy implies that $\mathcal{Z}_{\mathcal{P}}$ has nonempty Euclidean interior. The key for making the connection between consistency and the Euclidean interior is the following topological property.
@@
The next result shows the relevance of nondegeneracy for analysing the set $\mathcal{Z}_{\mathcal{P}}$ and the dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ for a system $F\in\mathbb{C}[p,{x}^{\pm}]^{r}$.
@@
The complex implicit function theorem [Huy05, Prop. 1.1.11] now gives that there exists an open Euclidean neighborhood $B$ of $p^{*}$ contained in $\widetilde{\mathcal{Z}}$ (the set $\mathcal{Z}$ for $\widetilde{F}$) and hence in $\mathcal{Z}$. Intersecting $B$ with $\mathcal{P}$, we obtain the first part of (i). By Lemma 2.11, $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense in $\mathbb{C}^{k}$.
For (ii), as the pair $(p^{*},x^{*})$ is a nondegenerate zero of $F$ and $\mathcal{I}$ is irreducible, Proposition 2.9(i) gives that $\mathcal{I}$ has the expected dimension. This fact, (i) and Theorem 2.2(i) give now the second part of (ii). ∎

### v4

@@ 1. Introduction
In a follow-up paper [FHPE24], we use the results from this paper to study the generic geometry of steady state varieties, and strengthen several previous statements about reaction networks, in particular concerning absolute concentration robustness from [PEF22, GPGH+25] and nondegenerate multistationarity from [CFMW17].
@@
We let $\star\colon\mathbb{C}^{n}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ denote the Hadamard product, given by $({x}\star{y})_{i}=x_{i}y_{i}$. For a field $K$, we let $K^{*}=K\setminus\{{0}\}$ be the group of units. For a matrix $A=(a_{ij})\in\mathbb{Z}^{n\times m}$ and a vector ${x}\in(K^{*})^{n}$, we let ${x}^{A}\in K^{m}$ be defined by $({x}^{A})_{j}=x_{1}^{a_{1j}}\cdots x_{n}^{a_{nj}}$ for $j=1,\ldots,m$. The zero matrix of size $n\times m$ is denoted $0_{n\times m}$, and ${\rm Id}_{n}$ is the identity matrix of size $n$. When saying that a property holds generically in a family indexed by some parameters in a set $\mathcal{P}\subseteq\mathbb{C}^{k}$, we mean that it holds in a nonempty open subset of $\mathcal{P}$, with respect to the subspace topology induced by the Zariski topology on $\mathbb{C}^{k}$. For a set $S\subseteq\mathbb{C}^{n}$, we let $\overline{S}$ denote the Zariski closure of $S$ in $\mathbb{C}^{n}$.
@@ 2.1. Framework
$\pi\colon\mathcal{I}\rightarrow\mathbb{C}^{k},\quad(p,{x})\mapsto p,$
where the system $F$ is implicit in the notation. For each choice of parameters $p\in\mathbb{C}^{k}$, we get a specialized system
@@
We say that $F$ is generically consistent if $\mathcal{Z}$ is Zariski dense in $\mathbb{C}^{k}$. Since $\mathcal{Z}$ is constructible (see, e.g., [CLO15, Theorem 3.2.3]), this is equivalent to $\mathcal{Z}$ containing a Zariski open subset of $\mathbb{C}^{k}$. Hence, if $F$ is generically consistent, a property holds generically in $\mathcal{Z}$ if and only if it holds generically in $\mathbb{C}^{k}$.
@@
Under mild assumptions, the choice of $\mathcal{P}$ does not affect generic consistency of the system.
Lemma 2.1.
Let $F\in\mathbb{C}[p,x^{\pm}]^{r}$ and $\mathcal{P}\subseteq\mathbb{C}^{k}$. Then the following holds:
  1. (i)
    If $\overline{\mathcal{Z}_{\mathcal{P}}}=\mathbb{C}^{k}$, then $\overline{\mathcal{Z}}=\mathbb{C}^{k}$.
  2. (ii)
    If $\overline{\mathcal{P}}=\mathbb{C}^{k}$, then the converse of (i) holds.
Proof.
Part (i) is immediate, since $\mathcal{Z}_{\mathcal{P}}\subseteq\mathcal{Z}$. For part (ii), note that if $\overline{\mathcal{Z}}=\mathbb{C}^{k}$, then $\mathcal{Z}$ contains a nonempty Zariski open subset $U$ of $\mathbb{C}^{k}$. The intersection $U\cap\mathcal{P}$ is Zariski dense in $\mathbb{C}^{k}$ since it is the intersection of a nonempty Zariski open and a Zariski dense set, and $\overline{\mathcal{Z}_{\mathcal{P}}}=\mathbb{C}^{k}$ follows. ∎
With this notation in place, we proceed now to study the generic dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ for $p\in\mathcal{P}$ in relation to whether the system is generically consistent. We then move on to study zeros in subsets $\mathcal{X}\subseteq(\mathbb{C}^{*})^{n}$. We also relate the study of nonemptiness and dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ to the concept of nondegeneracy. The connections among these properties are summarized in Theorem 2.17 towards the end of this section.
@@ 2.2. Generic consistency and dimension
Let $\mathcal{P}\subseteq\mathbb{C}^{k}$ and $F\in\mathbb{C}[p,{x}^{\pm}]^{r}$ be such that the incidence variety $\mathcal{I}$ is irreducible. Then:
@@
    For all $p\in\mathcal{Z}$, it holds that
    $\dim(Y)\geq\dim(\mathcal{I})-\dim(\overline{\mathcal{Z}})$   for all irreducible components $Y\subseteq\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$
    with equality for generic $p\in\overline{\mathcal{Z}}$.
@@
    If $\overline{\mathcal{Z}_{\mathcal{P}}}=\mathbb{C}^{k}$, then $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ has pure dimension $\dim(\mathcal{I})-k$ for generic $p\in\mathcal{Z}_{\mathcal{P}}$.
  3. (iii)
    If $\overline{\mathcal{Z}_{\mathcal{P}}}\subsetneq\mathbb{C}^{k}$ and $\overline{\mathcal{P}}=\mathbb{C}^{k}$, then all irreducible components of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ have dimension greater than $\dim(\mathcal{I})-k$ for all $p\in\mathcal{Z}$.
@@
Part (i) follows by applying the classical theorem of dimension of fibers (see, e.g., [Mum76, Theorem 3.13, Corollary 3.15]) to the canonical projection map
$\pi\colon\mathcal{I}\rightarrow\overline{\mathcal{Z}},\quad(p,x)\mapsto p.$
Part (ii) follows by noting that (i) and Lemma 2.1(i) together give that there is a nonempty Zariski open subset $U\subseteq\mathbb{C}^{k}$ such that $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ has pure dimension $\dim(\mathcal{I})-k$ for all $p\in U$. In particular, this holds for $p\in\mathcal{Z}_{\mathcal{P}}\cap U$, which is nonempty since $\overline{\mathcal{Z}_{\mathcal{P}}}=\mathbb{C}^{k}$. Finally, part (iii) follows directly from (i) and Lemma 2.1(ii). ∎
Combining part (ii) and (iii) of Theorem 2.3 when $\overline{\mathcal{P}}=\mathbb{C}^{k}$, we obtain that $\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))=\dim(\mathcal{I})-k$ for some $p\in\mathcal{Z}_{\mathcal{P}}$ if and only if $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense in $\mathbb{C}^{k}$.
@@
For the system in (1.8) with $n=k=r=2$, we have $\mathcal{Z}=(\mathbb{C}^{*})^{2}$ and $\mathcal{I}$ is irreducible of dimension $3$. By Theorem 2.3(ii), it holds that $\dim(\mathbb{V}_{\mathbb{C}^{*}}(F_{p}))=3-2=1$ for generic $p\in(\mathbb{C}^{*})^{2}$, which we already noticed in (1.8). In this case, the generic dimension differs from the expected dimension $n-r=0$ from (2.3).
We end by connecting consistency and dimension with the notion of flatness. Recall that a morphism of varieties $f\colon X\to Y$ is said to be flat at a point $x\in X$ if the induced homomorphism of local rings $\mathscr{O}_{Y,f(x)}\to\mathscr{O}_{X,x}$ is a flat ring homomorphism.
Proposition 2.5.
Let $F\in\mathbb{C}[p,x^{\pm}]^{r}$ be a parametric system such that $\mathcal{I}$ is irreducible. Then the following holds:
  1. (i)
    If the projection $\pi\colon\mathcal{I}\to\mathbb{C}^{k}$ is flat at generic points $(p,x)\in\mathcal{I}$, then $\overline{\mathcal{Z}}=\mathbb{C}^{k}$.
  2. (ii)
    If $\mathcal{I}$ is nonsingular, then the converse of (i) holds.
Proof.
It follows from [Sta24, 00ON] that if $\pi\colon\mathcal{I}\to\mathbb{C}^{k}$ is flat at a point $(p,x)\in\mathcal{I}$, then the local dimension of $\pi^{-1}(p)$ at $(p,x)$ is $\dim(\mathcal{I})-k$, and [Sta24, 00R4] gives that the converse is true if $\mathcal{I}$ is nonsingular (and hence in particular Cohen–Macaulay). The desired result now follows from Theorem 2.3. ∎
@@ 2.3. Restricting the ambient space
but equality might not hold (consider $F=(x_{1}-1)^{2}+(x_{2}-1)^{2}$ and $\mathcal{X}=\mathbb{R}^{2}_{>0}$). In general, equality in (2.4) holds if $\mathcal{X}$ is a Euclidean open subset of $(\mathbb{C}^{*})^{n}$, or if $\mathcal{X}$ is a Euclidean open subset of $(\mathbb{R}^{*})^{n}$ and additionally each irreducible component of $\mathbb{V}_{\mathbb{C}^{*}}^{\mathcal{X}}(F)$ contains a nonsingular real point (see [PEF22, Theorem 6.5] and [BCR98, Proposition 3.3.16]). Furthermore, if $F\in\mathbb{R}[x^{\pm}]^{r}$, then $n-r$ is a lower bound of the dimension of $\mathbb{V}_{\mathbb{R}^{*}}(F)$ if it is nonempty and each of its irreducible components contains a nonsingular point. Similarly, $n-r$ is a lower bound on the dimension of $\mathbb{V}_{\mathbb{R}^{*}}(F)\cap\mathbb{R}^{n}_{>0}$ as a semialgebraic set if it is nonempty and each irreducible component of $\mathbb{V}_{\mathbb{R}^{*}}(F)$ that intersects >$\mathbb{R}^{n}_{>0}$ contains a nonsingular point.
@@
If $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$ is Zariski dense in $\mathbb{C}^{k}$, then so is $\mathcal{Z}_{\mathcal{P}}$. If $\mathcal{I}$ is irreducible and attains its expected dimension $k+n-r$, then Theorem 2.3(ii) readily gives that $\mathbb{V}_{\mathbb{C}^{*}}^{\mathcal{X}}(F_{p})$ has pure dimension $n-r$ for generic $p\in\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$. The following example illustrates that Theorem 2.3(iii) might not extend to subsets $\mathcal{X}$ unless extra requirements are imposed.
@@
In particular $\mathcal{Z}$ is Zariski dense in $\mathbb{C}^{k}$ if and only if $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$ is.
@@
Assume that $\mathcal{I}$ is nonsingular and irreducible. If $\mathcal{P}$ and $\mathcal{X}$ are Euclidean open subsets of $\mathbb{C}^{k}$ and $(\mathbb{C}^{*})^{n}$, respectively, then $\mathcal{I}\cap(\mathcal{P}\times\mathcal{X})$ is Zariski dense in $\mathcal{I}$ if this intersection is nonempty. The same conclusion holds if $F$ has real coefficients and $\mathcal{P}$ and $\mathcal{X}$ are Euclidean open subsets of $\mathbb{R}^{k}$ and $\mathbb{R}^{n}$, respectively. This follows from the nonsingularity of $\mathcal{I}$, together with the fact that for a complex irreducible variety defined by polynomials with real coefficients that has at least one real nonsingular point, the real points form a Zariski dense subset of the complex variety [BCR98, Proposition 3.3.16]. Observe that in Example 2.7, the intersection $\mathcal{I}\cap(\mathcal{P}\times\mathcal{X})$, which is not dense in $\mathcal{I}$, consists only of singular points.
@@ 2.4. Nondegeneracy, Euclidean interior, and (real) dimension
Unlike the weaker notion of nonsingularity, nondegeneracy depends on the particular tuple $F\in\mathbb{C}[x^{\pm}]^{r}$, and not just the variety $\mathbb{V}_{\mathbb{C}^{*}}(F)$. The relationship between nondegenerate zeros and nonsingular points, and the connection to dimension, is summarized in the following proposition, which gathers several well-known results.
@@
Statement (i) is [CLO15, Theorem 9.6.9]. For (ii), the uniqueness follows from [CLO15, Theorem 9.6.8(iv)], whereas the density and dimensionality claims follow from Remark 2.6 (see also [BCR98, Theorem 3.3.10]). For statement (iii), consider the Jacobian $J_{F}=(\partial f_{i}/\partial x_{j})\in\mathbb{C}[p,x^{\pm}]^{r\times n}$, and note that the complement $\mathcal{I}\setminus\mathcal{U}\subseteq\mathcal{I}$ is cut out by the $r$-minors of $J_{F}$, and therefore forms a Zariski closed subset of $\mathcal{I}$. ∎
Theorem 2.3 allows us to determine the generic dimension of $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ from the Zariski denseness of $\mathcal{Z}_{\mathcal{P}}$, while the implicit function theorem (as we will use below in Proposition 2.14) gives that nondegeneracy implies that $\mathcal{Z}_{\mathcal{P}}$ has nonempty Euclidean interior. The key for making the connection between consistency and the Euclidean interior is the following topological property.
@@
The complex implicit function theorem [Huy05, Proposition 1.1.11] now gives that there exists an open Euclidean neighborhood $B$ of $p^{*}$ contained in $\widetilde{\mathcal{Z}}$ (the set $\mathcal{Z}$ for $\widetilde{F}$) and hence in $\mathcal{Z}$. Intersecting $B$ with $\mathcal{P}$, we obtain the first part of (i). By Lemma 2.13, $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense in $\mathbb{C}^{k}$.
For (ii), as the pair $(p^{*},x^{*})$ is a nondegenerate zero of $F$ and $\mathcal{I}$ is irreducible, Proposition 2.11(i) gives that $\mathcal{I}$ has the expected dimension. This fact, (i) and Theorem 2.3(ii) give now the second part of (ii). ∎